# WaveGNN: Modeling Irregular Multivariate Time Series for Accurate Predictions

### v1

@@ 5.1 Experimental Setup
Baselines. We compare WaveGNN to 8 methods: IP-Net [41], DGM2-O [30], mTAND [42] and UTDE [26] are imputation-based baselines, while GRU-D [29], SeFT [20], MTGNN [43], and Raindrop [19], are specialized models designed to handle irregularly sampled datasets. UTDE and Raindrop are considered the state-of-the-art imputation-based and specialized model approaches, respectively.
Table 2: Performance comparison with the state-of-the-art on P12, P19, and PAM datasets. P12 and P19 are evaluated on mortality and sepsis prediction, respectively, based on the AUROC and AUPRC metrics. The PAM dataset is evaluated on activity classification in terms of (weighted) Accuracy, Precision, Recall, and F1 Score. The reported values are averaged across 3 independent runs. Best and second best values are highlighted.
@@
WaveGNN 83.9 $\pm$ 1.2 49.4 $\pm$ 1.5 88.0 $\pm$ 0.9 57.1 $\pm$ 4.7 93.5 $\pm$ 1.1 93.7 $\pm$ 1.1 93.5 $\pm$ 1.1 93.5 $\pm$ 1.1
Improvement -0.59% +2.48% +1.14% +10.23% +3.42% +3.30% +3.42% +3.42%
Table 3: Performance comparison with the state-of-the-art on MIMIC III dataset, for 48-IHM and 24-PHE tasks. The performance of 48-IHM is measured on F1 and AUPRC, and 24-PHE on F1 (Macro) and AUROC. We report the average results across 3 independent runs. Best and second best values are highlighted.
@@
Improvement +2.64% -3.62% +23.69% -0.93%
5.2 Results
Overall Performance. We report the overall performance of WaveGNN and baselines in Tables 2 and 3, where the best and second best values are highlighted. WaveGNN consistently outperforms the baselines in the majority of cases (9 out of 12 experiments), and when it doesn’t, its performance is very close to the best-performing baseline.
Notably, WaveGNN achieves significant improvements in AUPRC, with gains of 10.2% and 2.4% over the best baseline in the binary tasks on the P19 and P12 datasets, respectively. In the multi-class task on the PAM dataset, WaveGNN consistently outperforms all baselines, demonstrating performance improvements of over 3.3% across all metrics. The PAM dataset has the smallest missing ratio and very long sequences among the evaluated datasets. This highlights that WaveGNN is highly effective not only in challenging scenarios like P19, which has over a 94% missing ratio and short sequences, but also in cases with less irregularity.
The results on the binary 48-IHM task of the MIMIC dataset are mixed. While WaveGNN outperforms UTDE by 2.64% in terms of F1 score, UTDE performs better in terms of AUPRC. This can be attributed to MIMIC’s longer sequences of observations with a lower missing data ratio compared to the P12 and P19 binary tasks, making imputation-based methods more effective for this less challenging binary classification task. However, in the more challenging multi-label classification task of 24-PHE on the same dataset, WaveGNN outperforms UTDE by a margin of 23.7% in F1 score, demonstrating that imputation-based methods become less effective in more complex scenarios.
In sum, on datasets with lower missing data ratios and longer sequences, like MIMIC and PAM, imputation-based methods like UTDE tend to outperform specialized models. Conversely, on datasets with higher missing data ratios, such as P19, specialized models like Raindrop are more effective. This makes sense: with more observations, imputation can more accurately estimate missing values, whereas in cases with fewer observations, the biased assumptions of imputation methods become less effective, making direct modeling approaches more suitable. Notably, WaveGNN consistently outperforms the best of both methods across all datasets.
Leave-sensors-out settings. We also evaluate WaveGNN under more challenging scenarios where whole sensor observations are missing. This simulates real-world cases where some sensors fail or are unavailable. Following the approach adopted by Raindrop [19], we experimented with two setups on PAM dataset. In the first one, we remove all the observations for a fixed set of the most informative sensors from samples, while in the second setting, we remove all the observations for a random set of sensors. Table 4 presents the results for both cases when 10%, 30%, and 50% of the sensors are removed. Our approach outperforms the baselines in 23 out of 24 experiments and ranks second in only one case. Specifically, WaveGNN demonstrates notable improvements, with up to 15.5% relative improvement in F1 score compared to the best baseline with 10% missing sensors, and up to 45.2% and 7.5% improvement for cases with 30% and 50% missing sensors, respectively.
Among the baselines, Raindrop remains the strongest competitor to WaveGNN. This shows the importance of capturing inter-series relationships and highlights the effectiveness of graph-based modeling. By successfully capturing the dependencies between sensors, missing data can be compensated, leveraging insights from correlated sensors that are relevant to the final outcome. In contrast, imputation-based methods, like mTAND and UTDE, demonstrate the largest performance degradation as the missing sensor ratio increases. These approaches still try to fill in the missing values, but with entire sensor observations missing, they have little to no reliable data to base their imputations on. This leads to inaccurate or biased reconstructions and, hence, affects their performance.
Table 4: Performance comparison on the PAM dataset under different missing sensor ratios and different settings. We report the average results across 3 independent runs. Best and second best values are highlighted.
@@ 5.2 Results
WaveGNN 43.2 $\pm$ 3.5 46.6 $\pm$ 6.1 43.2 $\pm$ 3.5 38.2 $\pm$ 3.4 51.2$\pm$ 3.3 63.3 $\pm$ 5.1 51.2 $\pm$ 3.3 51.2 $\pm$ 3.3
Improvement -7.29% +4.71% +1.88% +0.52% +3.01% +6.56% +14.28% +7.56%
5.3 Ablation Study
@@
We created different variations of WaveGNN to understand the effectiveness of its different components: w/o short-term and w/o long-term variants remove the $A_{S}$ and $A_{L}$ similarity matrices, so that the short- and long-term inter-series relationships are not considered, respectively. The w/o inter-series variant excludes the graph message passing part, so neither of the inter-series similarity measures are considered, while the w/o intra-series variant removes the transformer-based module and simply aggregates the observation embeddings for each sensor to obtain a univariate sequence embedding. Finally, the w/o temporal encoding variant discards the temporal encoding module and uses a simple positional encoding for the Transformer, and the w/o decay rate variant removes WaveGNN’s decay mechanism, aggregating the output sequence embedding of the Transformer with equal weights, thus treating all observations in a sequence equally.
@@
Table 5 presents the results of the ablation study on the PAM dataset, showing the impact of removing various components of WaveGNN. Our findings indicate that all components contribute to improved performance, with the most significant drops occurring when inter-series and intra-series relationships are removed, leading to performance decreases of 33.8% and 13.2% in weighted F1 score, respectively. This shows the importance of simultaneously addressing both intra-variable dynamics and inter-variable dependencies to effectively model irregular datasets.
@@
Table 5: Ablation study results on the PAM dataset. We report the average results across 3 independent runs.
@@
WaveGNN 93.5 $\pm$ 1.1 93.7 $\pm$ 1.1 93.5 $\pm$ 1.1 93.5 $\pm$ 1.1
6 Conclusion
In this study, we introduced WaveGNN, a novel framework designed to directly embed irregularly sampled multivariate time series datasets for accurate predictions. Unlike previous imputation-based and specialized model approaches, WaveGNN simultaneously captures both intra-series correlations and inter-series relationships, enabling the model to effectively compensate for missing observations. Our evaluation demonstrated that WaveGNN outperforms state-of-the-art baselines across diverse datasets, even under conditions of extreme sparsity where entire sensors are missing, thereby validating the importance of each proposed component within the framework. In the future, we plan to expand WaveGNN to support multimodal scenarios.

### v2

@@ V-A Experimental Setup
Baselines. We compare WaveGNN to 11 methods. IP-Net [4], DGM2-O [12], mTAND [5], and UTDE [13] are interpolation-based models. GRU-D [6], SeFT [2], MTGNN [27], and WarpFormer [10] are specialized architectures for modeling irregular time series. Raindrop [1], T-PatchGNN [9], and GraFITi [11] are graph-based methods that model inter-variable relationships under irregular sampling. Among these, GraFITi and WarpFormer represent recent state-of-the-art models.
Dataset Preprocessing. To ensure compatibility with previous approaches, we followed the preprocessing steps outlined in [26]  [^2] https://github.com/YerevaNN/mimic3-benchmarks/tree/master/mimic3benchmark for the MIMIC III dataset. For the P12, P19, and PAM datasets, we used the preprocessed data provided by [1]  [^3] https://figshare.com/articles/dataset/P19_dataset_for_Raindrop/19514338/1?file=34683070444https://figshare.com/articles/dataset/P12_dataset_for_Raindrop/19514341/1?file=34683085555https://figshare.com/articles/dataset/PAM_dataset_for_Raindrop/19514347/1?file=34683103. To compare our approach with baselines, we used the same data split indexes as provided by [1] for P12, P19, and PAM, and by [26] for MIMIC.
Computational Resources Our experiments were performed on a cluster node equipped with a 24 GB NVIDIA GeForce RTX 3090 GPU with CUDA version 11.8 and a Ryzen 9 5900 12-core processor, running on Ubuntu 20.04. Furthermore, all neural network models are implemented based on PyTorch version 2.2.2 and Pytorch Geometric version 2.5.3 on Python version 3.9.19.
Hyperparameter Configuration Training Setup. All experiments were run for 20 epochs using AdamW [28] with a learning rate of 1e-4, dropout rate of 0.1, and weight decay of 1e-5. A batch size of 4 was used with gradient accumulation over 32 steps to achieve an effective batch size of 128. We employed a ReduceLROnPlateau scheduler (factor 0.1, patience 1, min LR 1e-8). Model selection was based on AUROC (binary tasks), weighted F1 (PAM), and AUPRC (24-PHE), with early stopping (patience 10). Each experiment was repeated 3 times, and mean and standard deviation were reported.
Input Window. Full observation windows were used for P12 (max 215), P19 (60), and PAM (600). For MIMIC, we used the last 700 timestamps to ensure memory efficiency.
WaveGNN Configuration. Static features (when available) were embedded via a 2-layer MLP with ReLU activations to $Z_{i}\in\mathbb{R}^{128}$. Observation embeddings $h_{i,v}^{t}\in\mathbb{R}^{64}$ were generated via a 3-layer MLP with $\tanh$ activations. Sensor sequences were encoded using a 2-layer Transformer with 64-dim outputs, using 64-dim relative time encodings and an initial decay parameter $\eta=0.1$. A linear projection then mapped outputs to 128-dim.
Short-term sensor similarity ($A_{S}$) was computed using 8-head self-attention. Long-term context came from a learnable global embedding matrix $\tilde{Z}\in\mathbb{R}^{\#sensors\times 128}$; $\alpha$ was initialized to 0.5 to balance $A_{S}$ and $A_{L}$. To maintain sparsity, we pruned the lowest 30% of edges.
The GNN module consisted of 3 GCN layers with residual connections, layer norm, and output dimension 128. The graph embedding module concatenated (1) a 3-layer MLP-transformed global node embedding and (2) a feature-wise max pooling over all nodes, yielding a 256-dim vector. This was passed through a final 2-layer MLP with ReLU to produce the task-specific output.
Additional hyperparameter details are provided in the code readme.
TABLE II: Performance comparison with the state-of-the-art on P12, P19, and PAM datasets. P12 and P19 are evaluated on mortality and sepsis prediction, respectively, based on the AUROC and AUPRC metrics. The PAM dataset is evaluated on activity classification in terms of (weighted) Accuracy, Precision, Recall, and F1 Score. The reported values are averaged across 3 independent runs. Best and second best values are highlighted.
@@
T-PatchGNN 62.1 $\pm$ 4.8 21.3 $\pm$ 1.8 82.0 $\pm$ 3.2 52.1 $\pm$ 2.3 90.6 $\pm$ 1.6 90.9 $\pm$ 1.8 90.6 $\pm$ 1.6 90.6 $\pm$ 1.7
Warpformer 81.9 $\pm$ 2.8 39.6 $\pm$ 1.6 89.4 $\pm$ 0.3 64.9 $\pm$ 0.4 91.4 $\pm$ 1.4 91.5 $\pm$ 1.3 91.4 $\pm$ 1.4 91.4 $\pm$ 1.4
GraFITi 80.3 $\pm$ 1.0 37.7 $\pm$ 0.4 90.5 $\pm$ 1.1 66.0 $\pm$ 1.7 94.6 $\pm$ 0.5 94.7 $\pm$ 0.5 94.6 $\pm$ 0.5 94.6 $\pm$ 0.5
WaveGNN 83.9 $\pm$ 1.2 49.4 $\pm$ 1.5 88.0 $\pm$ 0.9 57.1 $\pm$ 4.7 95.6 $\pm$ 1.1 95.7 $\pm$ 1.1 95.6 $\pm$ 1.1 95.6 $\pm$ 1.1
TABLE III: Performance comparison with the state-of-the-art on MIMIC III dataset, for 48-IHM and 24-PHE tasks. The performance of 48-IHM is measured on F1 and AUPRC, and 24-PHE on F1 (Macro) and AUROC. We report the average results across 3 independent runs, highlighting best and second best values.
@@
TABLE IV: Cross-dataset/task consistency across all 12 dataset-metric evaluation pairs (P12, P19, PAM, and MIMIC). We report average rank (lower is better), the number of times each method ranked first (Top-1), and the number of times each method ranked among the top two (Top-2).
Method Avg Rank Top-1 Count Top-2 Count
WaveGNN 1.67 7 9
UTDE 4.58 2 4
GraFITi 5.25 2 6
Warpformer 5.04 0 2
Raindrop 5.50 0 0
DGM2-O 6.71 1 1
GRU-D 6.62 0 0
mTAND 6.88 0 2
T-PatchGNN 7.25 0 0
IP-Net 7.88 0 0
MTGNN 8.79 0 0
SeFT 11.83 0 0
TABLE V: Robustness experiments on the PAM dataset under different missing sensor ratios and different settings. We report the average results across 3 independent runs. Best and second best values are highlighted.
@@
GraFITi 22.9 $\pm$ 1.2 10.6 $\pm$ 0.9 22.9 $\pm$ 1.2 12.4 $\pm$ 1.9 47.9 $\pm$ 3.4 53.4 $\pm$ 4.4 47.9 $\pm$ 3.4 48.4 $\pm$ 4.5
Warpformer 43.6 $\pm$ 3.2 35.3 $\pm$ 3.4 43.6 $\pm$ 3.2 35.0 $\pm$ 2.7 53.6 $\pm$ 2.1 54.0 $\pm$ 2.4 53.6 $\pm$ 2.1 50.7 $\pm$ 3.4
T-PatchGNN 31.0 $\pm$ 1.7 15.6 $\pm$ 6.0 31.0 $\pm$ 1.7 20.0 $\pm$ 5.1 40.1 $\pm$ 1.8 29.5 $\pm$ 6.1 40.1 $\pm$ 1.8 31.6 $\pm$ 0.8
[…]
Improvement -7.29% +4.71% -0.91% +0.52% -4.47% +6.56% -4.47% +0.99%
V-B Results
Overall Performance. We report the overall performance of WaveGNN and competing baselines in Tables II and III, highlighting the best and second best scores.
WaveGNN achieves the best performance on PAM across all metrics, underscoring its strength in multi-class activity classification under high sparsity. On P12, it provides the top AUPRC (49.4) and an AUROC nearly tied with DGM2-O (83.9 vs. 84.4, a 0.5 gap), showing competitive performance on highly irregular mortality prediction. On P19, GraFITi achieves higher scores, because this binary task involves much shorter sequences, favoring interpolation-based methods.
As shown in Table III, on the MIMIC dataset, WaveGNN shows mixed results on the binary 48-IHM task. While WaveGNN improves over the next best model, UTDE, by 2.64% in F1 score, UTDE achieves a higher AUPRC. This can be attributed to MIMIC’s lower missingness compared to P12 and P19, which makes interpolation-based models more competitive in this relatively easier binary task. However, in the more complex multi-label classification task of 24-PHE on the same dataset, WaveGNN outperforms UTDE by 23.7% in F1 score, suggesting that imputation-based approaches are less effective in challenging, multi-label settings.
Among graph-based baselines, Raindrop and GraFITi achieve competitive results in specific tasks (e.g., GraFITi on P19), but their performance fluctuates sharply across datasets. Conversely, WaveGNN consistently ranks among the top methods under varying levels of irregularity and sequence lengths.
Consistency Across Datasets and Tasks. To evaluate consistency in performance beyond individual benchmarks, we compute the average rank of each method across all 12 dataset-metric pairs. Table IV shows the results. WaveGNN achieves the lowest average rank (1.67), substantially ahead of the next best baseline (UTDE at 4.58). Moreover, WaveGNN is ranked first in 7 cases and within the top two in 9 cases, whereas competing methods rarely maintain such consistency (e.g., UTDE achieves top performance on some MIMIC evaluation metrics but ranks poorly on P12 and PAM, GraFITi excels on P19 but drops sharply on P12). These results demonstrate that WaveGNN provides consistent performance across datasets and tasks, avoiding the inconsistency of baselines that perform well only under favorable conditions.
Robustness to Missing Data. We evaluate WaveGNN under more challenging conditions where entire sensor streams are missing, simulating real-world scenarios such as sensor failures or deployment constraints. Following Raindrop [1], we experiment with two setups on the PAM dataset: (1) removing all observations from a fixed set of the most informative sensors, and (2) removing observations from a randomly selected set of sensors. Table V reports the results for both settings when 10%, 30%, and 50% of sensors are removed.
WaveGNN outperforms all baselines in 20 out of 24 experiments and ranks second in the remaining four. Notably, it achieves up to 5% relative improvement in F1 score with 10% sensor removal, and gains of 12.3% and 6.5% with 30% and 50% removal, respectively.
Among the baselines, Raindrop and WarpFormer are the most competitive in robustness. In contrast, GraFITi, which previously ranked among the top performers, suffers a significant performance drop. This drop is likely due to GraFITi’s reliance on inter-variable interactions at shared timestamps. When informative sensor observations are missing, the model has less reliable information to compensate for irregularity.
Imputation-based methods such as mTAND and UTDE also experience substantial declines in robustness with increasing sensor removal. These approaches still try to fill in the missing values, but with entire sensor observations missing, they have little to no reliable data to base their imputations on. This leads to inaccurate or biased reconstructions which affects their robustness under missing data.
V-C Ablation Study
@@
We created different variations of WaveGNN to understand the effectiveness of its components: w/o short-term and w/o long-term variants remove the $A_{S}$ and $A_{L}$ similarity matrices, so that the short- and long-term inter-series relationships are not considered, respectively. The w/o inter-series variant excludes the graph message passing part, so neither of the inter-series similarity measures are considered, while the w/o intra-series variant removes the transformer-based module and simply averages the observation embeddings for each sensor to obtain a univariate sequence embedding. Finally, the w/o temporal encoding variant discards the temporal encoding module and uses a simple positional encoding for the Transformer, and the w/o decay rate variant removes WaveGNN’s decay mechanism, aggregating the output sequence embedding of the Transformer with equal weights, thereby treating all sequence observations equally.
@@
Table VI presents the results of the ablation study on the PAM dataset, showing the impact of removing various components of WaveGNN. Our findings indicate that all components contribute to improved performance, with the most significant drops occurring when inter-series and intra-series relationships are removed, leading to performance decreases of 35.3% and 15.1% in weighted F1 score, respectively. This shows the importance of simultaneously addressing both intra-variable dynamics and inter-variable dependencies to effectively model irregular datasets.
@@
(a) Dynamic Short-Term Dependencies ($A_{S}$) - Sample 1
(b) Dynamic Short-Term Dependencies ($A_{S}$) - Sample 2
(c) Static Long-Term Similarities ($A_{L}$)
Figure 2: Learned adjacency matrices capturing feature dependencies in ICU patients. (a) and (b) show dynamic short-term dependencies ($A_{S}$) for two test samples, highlighting patient-specific variations. (c) presents the static long-term similarity matrix ($A_{L}$), representing global feature relationships shared across all samples.
TABLE VI: Ablation study results on the PAM dataset. We report the average results across 3 independent runs.
@@
WaveGNN 95.6 $\pm$ 1.1 95.7 $\pm$ 1.1 95.6 $\pm$ 1.1 95.6 $\pm$ 1.1
V-D Interpretation of Learned Sensor Graphs
In this section, we examine the interpretability of the sensor graphs constructed by WaveGNN, using a trained model checkpoint on the P12 dataset. Specifically, we analyze the adjacency matrices learned for two distinct test samples, which were excluded from the training set. The first sample corresponds to a patient who remained in the ICU for more than three days ($\text{label}=1$), while the second corresponds to a patient discharged within three days ($\text{label}=0$).
As detailed in Section IV, WaveGNN constructs two types of adjacency matrices to model graph-based relationships: a static, globally shared matrix $A_{L}$, encoding long-term similarities, and dynamic, sample-specific matrix $A_{S}$, encoding short-term dependencies within the input window. These matrices capture directed pairwise relationships between sensor features, where $A_{i,j}$ represents the influence of feature $i$ on feature $j$. The final adjacency matrix is derived as a weighted combination of $A_{L}$ and $A_{S}$, controlled by a trainable parameter $\alpha$ (Equation 13).
Figure 2 presents heatmaps of the learned adjacency matrices. Rows and columns represent sensor features, while the color intensity depicts the edge weight, indicating the strength of pairwise feature relationships.
Static Long-Term Similarities ($A_{L}$)
The globally shared matrix $A_{L}$ (Figure 2c) captures consistent long-term feature dependencies across the dataset. For example, FiO2 (fraction of inspired oxygen) and SaO2 (oxygen saturation) exhibit a relatively strong relationship (highlighted with red circles), aligned with the clinical understanding that FiO2 directly impacts blood oxygen levels [29]. This global dependency suggests that WaveGNN effectively captures universal feature relationships important for modeling patient states.
Dynamic Short-Term Dependencies ($A_{S}$)
The dynamic matrices $A_{S}$ (Figures 2a and 2b) are generated for each sample by WaveGNN’s learned graph constructor, adjusting the graph to each patient’s unique characteristics. These matrices encode short-term dependencies that vary across patients:
  • For the sample with $\text{label}=1$ (Figure 2a), Glasgow Coma Scale (GCS) and Cholesterol exhibit strong connections to other features (highlighted with red bars), suggesting their importance in predicting prolonged ICU stays. GCS is widely used to assess neurological status, and lower scores are often associated with critical conditions requiring extended ICU care [30]. Cholesterol levels are linked to metabolic and inflammatory responses, which can influence recovery time, with abnormalities often observed in patients with prolonged hospitalization [31]. The model’s identification of these relationships aligns with clinical expectations, indicating that WaveGNN effectively captures key physiological dependencies associated with extended ICU stays.
  • For the sample with $\text{label}=0$ (Figure 2b), Blood Urea Nitrogen (BUN) and Heart Rate (HR) exhibit strong connections to other features (highlighted with red bars), suggesting their role in characterizing shorter ICU stays. BUN is commonly used to monitor kidney function and metabolic stress, while HR is a key vital sign reflecting physiological stability. The learned graph structure indicates that these features are highly interconnected in short-stay patients, aligning with the idea that acute physiological changes tend to be more immediate and short-lived [32]. In contrast, prolonged ICU stays often involve more complex dependencies beyond acute variations, as seen in the previous case with GCS and Cholesterol. These differences in feature interactions for the two samples suggest that WaveGNN effectively captures short-term physiological variations, adjusting its graph structure to reflect patient-specific ICU trajectories.
Notably, the values in the short-term dependency matrices ($A_{S}$) are smaller compared to the long-term similarity matrix ($A_{L}$), as $A_{S}$ refines the graph structure for each sample rather than encoding global dependencies. The balance between these components is controlled by the learned parameter $\alpha$, which determines the relative contributions of $A_{L}$ and $A_{S}$ in the final graph. For this trained checkpoint, the model has learned $\alpha=0.5211$, indicating that short-term dependencies ($A_{S}$) contribute slightly more than long-term similarities ($A_{L}$).
VI Conclusion
We introduced WaveGNN, a novel framework for modeling irregular multivariate time series that avoids imputation and operates directly on raw clinical data. By combining a decay-aware Transformer for intra-series encoding with a sample-specific graph neural network for inter-series modeling, WaveGNN produces sparse, interpretable graphs per instance. Our evaluation shows that WaveGNN achieves consistent performance across four clinical benchmarks, obtaining the lowest average rank across 12 evaluations. The model is also robust to varying levels of missingness, outperforming all baselines in 20 of 24 missing-data robustness experiments, while providing interpretable graph structures that align with known physiological relationships. In the future, we plan to expand WaveGNN to support multimodal scenarios.